# SPECIAL RELATIVITY

### v11

@@ 5.13 Ether Theory and the Aberration of Light Effect
We begin by examining the influence of the ether wind on light speed. According to the ether hypothesis, the velocity of light observed from an accelerated frame would be $c+u$ for a beam propagating in the same direction as the accelerated system and $-c+u$ for a beam propagating in the opposite direction, where $u=-v$ represents the ether velocity in the accelerated reference frame. The aberration effect arises as a consequence of the ether wind’s influence on the speed of light (Fig. 23).
@@ 5.13.1 Lorentz’s Theory of Corresponding States
In the preceding discussion, we analyzed the aberration of light in an accelerated frame using the Fourier transform method. The fundamental asymmetry arises from the electrodynamics equation (Eq. 17). One consequence of the cross-term in the wave equation is the emergence of group velocity, which contributes to the aberration. However, an alternative approach provides an equally satisfactory explanation. This involves clock resynchronization, wherein Einstein’s synchronization method is used to define a Lorentz coordinate system for the optical setup at rest in the accelerated frame. In this framework, Maxwell’s equations describe the field configuration, and the resynchronized time $t^{\prime}_{n}$ is given by $t^{\prime}_{n}=t_{n}-x_{n}v/c^{2}$ in the first-order approximation. After resynchronization, the asymmetry shifts to the initial condition: the plane wavefront of the emitter undergoes a rotation by an angle $-v/c$ in the accelerated frame.
@@ 5.13.2 Analogy with the Sound Theory
where coordinates and time are transformed according to the Galilean transformation $x_{n}=x-vt$, $t_{n}=t$. To describe sound propagation, the wave equation must be integrated with the initial conditions of the sound wavefront. Notably, acceleration does not influence the wavefront’s orientation, meaning the emitted sound beam remains perpendicular to the $z_{n}$ axis. Since the air remains stationary in the atmospheric frame, any reference frame moving relative to it experiences anisotropy in sound propagation. As a result, an observer in such a moving frame perceives an ”air wind” when measuring the speed of sound. According to the wave equation (Eq. 19), the sound velocity observed from an accelerated reference frame is $v_{s}-v$ for a beam propagating in the same direction as the accelerated system and $-v_{s}-v$ for a beam propagating in the opposite direction. Here, $-v$ represents the wind velocity in the accelerated frame. This effect, known as aberration, arises due to the influence of the perceived air wind on the speed of sound (Fig. 23).
@@ 6.3 Stellar aberration in the Context of Special Relativity
It is widely accepted that the theory of relativity aligns with Bradley’s discovery of stellar aberration, interpreting it as a consequence of an observer’s motion relative to a light source 232323 In his paper on relativity, Einstein derived the aberration formula by considering the velocity $v$ as the relative velocity of the star-Earth system. This interpretation has been adopted by numerous textbook authors. As Moeller [13] states: ”This phenomenon, known as aberration, was observed by Bradley, who noticed that stars appear to undergo an annual collective motion in the sky. This apparent motion arises because the observed direction of a light ray from a star depends on the Earth’s velocity relative to the star.”.
@@ 7.1 Inertial frame view of Observations of a Non-Inertial Observer
In this chapter, we introduce a novel approach to the theory of light aberration in non-inertial reference frames, providing an alternative perspective on this complex issue. First, we highlight a fundamental principle: the laws of physics in initial inertial frame must account for all physical phenomena, including observations made by non-inertial observers. It is commonly believed that phenomena in non-inertial (e.g., rotating) reference frames should be analyzed solely within the geometric space-time framework using the Langevin metric. However, we argue that physical phenomena in non-inertial frames can also be effectively described within an inertial frame using standard Einsteinian special relativity, leveraging well-established relativistic kinematic effects.
For example, when observed from an inertial laboratory frame, the interpretation of the Sagnac effect is straightforward. The phase difference between counter-propagating waves, attributed to the Sagnac effect, can be derived from the relativistic velocity addition formula. In other words, within the rotating frame, as seen from the inertial lab frame, the Sagnac effect manifests as a purely kinematic phenomenon governed by Einstein’s velocity addition in special relativity. This analysis assumes that the speed of light in the lab frame is $c$, which, according to special relativity, remains independent of the motion of the source. In contrast, calculating the Sagnac effect within a rotating reference frame is more complex. In such cases, textbook authors often employ a metric tensor—specifically, the Langevin metric—within a four-dimensional Minkowski spacetime to determine the propagation time difference between counter-propagating waves. Could we not analyze the aberration of light in the same manner? Here, too, we apply the Langevin metric in a non-inertial reference frame.
@@
It is particularly interesting to observe that geometric effects in our ordinary space are closely linked to the relativity of simultaneity. In Fig. 16, the transmitted light beam propagates at an angle of $-v/c$, resulting in the phenomenon of light aberration. This naturally raises the question: relative to what does a light beam propagate within an accelerated frame when it undergoes an angular displacement of $-v/c$? Consider an observer in the accelerated frame measuring the direction of the light beam. If the plane wavefront of the transmitted beam is focused by a lens, it will form a diffraction spot in the focal plane along the optical axis. Determining the direction of the optical axis with respect to the frame’s axes is equivalent to measuring the angular displacement. To detect the aberration of light within the accelerated frame, it is essential to establish a coordinate system with a well-defined reference direction. This necessitates a detailed examination of the method by which coordinates are assigned—a process that inherently involves a physical procedure. Ultimately, this method must provide a consistent framework for defining coordinates in both inertial and accelerated reference frames.
@@
Now, let us return to the observations of an accelerated observer as viewed from an inertial frame. We begin with the non-covariant approach. The aberration of light problem can be addressed within a ”single inertial frame” description, without invoking Lorentz transformations. According to our approach, when a transversely moving screen with a hole is placed in the path of a normally incident plane wave of light, the transmitted light undergoes a deviation in energy transport (see Fig. 9). However, an alternative and equally valid explanation for the aberration of light from a moving source relies on Lorentz coordinatization. This explanation describes the uniform translational motion of the light source within the inertial frame.
In Chapter 3, we discussed how absolute time coordinatization can be transformed into Lorentz coordinatization. Within the inertial frame, this transformation can be understood as a change of the time variable according to $t\to t+xv/c^{2}$. The combined effect of the Galilean transformation and this variable change results in the Lorentz transformation when using absolute time coordinatization in the inertial frame. Consequently, in this framework, the speed of light remains independent of the motion of the source.
@@
From the preceding discussion, it follows that in an accelerated system, as viewed by an inertial observer, the electromagnetic axis assigned by the accelerated observer will not be parallel to the $z$-axis of the inertial frame (see Fig. 30). We conclude that, in the accelerated frame, the transmitted light beam propagates along the $z_{n}$-axis of the accelerated frame with an angular displacement of $-v/c$.
@@
Now, we examine the angular displacement of the transmitted light relative to the reference beam within the accelerated frame. On one hand, the angular displacement of the transmitted light with respect to the reference beam from an emitter at rest in the accelerated frame is $-v/c$. On the other hand, the angular displacement of the inertial reference light beam is $-2v/c$ (see Fig. 17). This implies that the accelerated observer can directly measure the angular displacement of the transmitted light with respect to the inertial frame’s reference light beam. Specifically, the relative angular displacement is $\theta_{a}(\mathrm{transmitted~{}beam})-\theta_{a}(\mathrm{inertial~{}reference~{}beam})=-v/c+2v/c=v/c$. This result is consistent with the measurements of the inertial observer, as expected  [^26] We have been discussing the electromagnetic axis. If we instead determine the reference directions in both frames using two relativistic electron sources, an important distinction arises. Unlike a reference light beam, a reference relativistic electron beam would not allow for an exact measurement of the transmitted light’s angular displacement within our experimental setup..
@@ 7.4 Discussion
For completeness, we must also analyze the angular displacement of the $x_{n}$-axis. So far, we have considered only non-relativistic speeds, employing a first-order relativistic approximation. It follows that for the displacement of the detector along the $x$-axis in the $z$-direction, we obtain: $\Delta z\simeq\theta_{w}Lv/c\simeq Lv^{2}/c^{2}$. Thus, special relativity indicates that in the limit $v^{2}/c^{2}\to 0$, the kinematic shearing motion (during acceleration) and the Wigner rotation motion become indistinguishable.
@@ 8 Kinematics of Wigner Rotation
The focus of this chapter is the Wigner rotation  [^28] It is well known that the composition of non-collinear Lorentz boosts does not result in a simple boost but rather in a Lorentz transformation that includes both a boost and a spatial rotation. This rotation is known as the Wigner rotation [39, 40]. It is sometimes referred to as the Thomas rotation (see, e.g., [13, 41]).. We previously introduced the Wigner rotation in the last chapter, where our discussion primarily centered on the aberration of light. However, the Wigner rotation is also closely related to the aberration of particles. As a relativistic kinematic effect, the Wigner rotation describes how the coordinate axes of a reference frame, moving along a curvilinear trajectory, rotate with respect to the axes of an inertial Lorentz frame. The primary objective of this chapter is to derive an expression for this effect. Throughout, we have aimed to minimize mathematical complexity while maintaining clarity.
@@ 8.1 Composition of Lorentz Boosts
Let us denote the three inertial frames by $K,K^{\prime},K"$. The initial inertial frame is $K$, while $K^{\prime}$ is the rest frame of the particle, which has a velocity $\vec{v}=\vec{v}(\tau)$ relative to $K$, and $K"$ represents the particle’s rest frame at the next instant of proper time $\tau+d\tau$, moving with an infinitesimal velocity $d\vec{v^{\prime}}$ relative to $K^{\prime}$. All three inertial frames are assumed to be Lorentz frames. To clarify velocity measurements in different inertial frames: an observer in $K$ measures the velocity of frame $K^{\prime}$ as $\vec{v}$, while an observer in $K^{\prime}$ measures the velocity of frame $K"$ as $d\vec{v^{\prime}}$.
Now, let us examine the transformation of velocity in the theory of relativity. For a rectilinear motion along the $z$ axis, the velocity transformation follows the equation: $v_{z}(\tau+d\tau)=(dv^{\prime}_{z}+v_{z})/(1+v_{z}dv^{\prime}_{z}/c^{2})$. As it happens with the composition of Galilean boosts, collinear Lorentz boosts commute. This implies that the result of successive collinear Lorentz boosts is independent of which transformation applies first.
Unlike the case of Lorentz boosts in collinear directions, Lorentz boosts in different directions do not commute. While the successive application of two Galilean boosts results in a single Galilean boost, and the successive application of two rotations yields another rotation, the composition of two non-collinear Lorentz boosts does not produce a single Lorentz boost.
We assume that frame $K^{\prime}$ moves with velocity $v_{z}$ along the $x$-axis relative to $K$ and in frame $K^{\prime}$, the particle’s acceleration is perpendicular to the trajectory of the initial inertial frame. Suppose that frame $K"$ moves relative to $K^{\prime}$ with nonrelativistic velocity $dv^{\prime}_{x}$.
From the perspective of an inertial observer, time dilation occurs in the Lorentz frame moving with velocity $\vec{v}_{z}$ relative to the initial inertial frame, given by $dt/\gamma=d\tau$. Due to this time dilation, the velocity increment in the initial inertial frame $dv_{x}$ corresponds to a velocity $dv^{\prime}_{x}=\gamma dv_{x}$ in frames $K^{\prime}$.
@@ 8.2 Wigner Rotation
[…]
@@
Transforming this expression back to the lab frame $K$ yields $d\vec{\Phi}_{K}=(1-1/\gamma)\vec{v}\times d\vec{v}/v^{2}$ representing the Wigner rotation angle of the proper frame’s axes as observed in the lab frame. According to the principles of special relativity, this quantity must remain form-invariant under Lorentz transformations  [^29] The correct expression for the Thomas (Wigner) rotation was first obtained by V. Ritus [42]. In deriving expressions for the Thomas (Wigner) rotation, the majority of authors (see e.g. [41]) was supposedly guided by the incorrect expression for Thomas (Wigner) rotation from Moeller’s monograph [13]. The expression obtained by Moeller is given by $\vec{\delta\Phi}=(1-\gamma)\vec{v}\times d\vec{v}/v^{2}=(1-\gamma)\vec{\delta\theta}$ (and subsequently expression for Thomas precession $\Omega_{\mathrm{T}}=(1-\gamma)\omega_{0}$). It should be noted that, in his monograph, Moeller stated several times that this expression is valid in the lab Lorentz frame. Clearly, this expression and correct result differ both in sign and magnitude. An analysis of why Moeller obtained an incorrect expression for the Thomas (Wigner) rotation in the lab frame is the focus of Ritus paper [43]. As analyzed in Ritus’s work [43], Moeller’s error is not computational but rather conceptual. The correct expression, derived decades earlier by several researchers, remained largely unnoticed amid the proliferation of inaccurate results, as discussed in the review [44].. We note that owing to the relativistic effect of time dilation in the reference frame that moves to the lab frame, the Wigner rotation angle in the proper frame is always $\gamma$ time higher than in the lab frame  [^30] We note that in 1986, M. Stranberg obtained an expression for the Thomas (Wigner) rotation correct both in the lab inertial frame and the comoving reference frame [45]. His paper is one of the few that explicitly states that the angle of Thomas (Wigner) rotation in the comoving frame is $\gamma$ times higher than in the lab frame..

### v12

@@ 5.13 Ether Theory and the Aberration of Light Effect
We begin by examining the influence of the ether wind on light speed. According to the ether hypothesis, the velocity of light observed from an accelerated frame would be $c+u$ for a beam propagating in the same direction as the accelerated system and c+u-c+u for a beam propagating in the opposite direction, where u=vu=-v represents the ether velocity in the accelerated reference frame. The aberration effect arises as a consequence of the ether wind’s influence on the speed of light (Fig. 23).
@@ 5.13.1 Lorentz’s Theory of Corresponding States
In the preceding discussion, we analyzed the aberration of light in an accelerated frame using the Fourier transform method. The fundamental asymmetry arises from the electrodynamics equation (Eq. 17). One consequence of the cross-term in the wave equation is the emergence of group velocity, which contributes to the aberration. However, an alternative approach provides an equally satisfactory explanation. This involves clock resynchronization, wherein Einstein’s synchronization method is used to define a Lorentz coordinate system for the optical setup at rest in the accelerated frame. In this framework, Maxwell’s equations describe the field configuration, and the resynchronized time $t^{\prime}_{n}$ is given by $t^{\prime}_{n}=t_{n}-x_{n}v/c^{2}$ in the first-order approximation. After resynchronization, the asymmetry shifts to the initial condition: the plane wavefront of the emitter undergoes a rotation by an angle v/c-v/c in the accelerated frame.
@@ 5.13.2 Analogy with the Sound Theory
where coordinates and time are transformed according to the Galilean transformation $x_{n}=x-vt$, $t_{n}=t$. To describe sound propagation, the wave equation must be integrated with the initial conditions of the sound wavefront. Notably, acceleration does not influence the wavefront’s orientation, meaning the emitted sound beam remains perpendicular to the $z_{n}$ axis. Since the air remains stationary in the atmospheric frame, any reference frame moving relative to it experiences anisotropy in sound propagation. As a result, an observer in such a moving frame perceives an ”air wind” when measuring the speed of sound. According to the wave equation (Eq. 19), the sound velocity observed from an accelerated reference frame is $v_{s}-v$ for a beam propagating in the same direction as the accelerated system and $v_{s}-v$ for a beam propagating in the opposite direction. Here, v-v represents the wind velocity in the accelerated frame. This effect, known as aberration, arises due to the influence of the perceived air wind on the speed of sound (Fig. 23).
@@ 6.3 Stellar aberration in the Context of Special Relativity
It is widely accepted that the theory of relativity aligns with Bradley’s discovery of stellar aberration, interpreting it as a consequence of an observer’s motion relative to a light source. 232323 In his paper on relativity, Einstein derived the aberration formula by considering the velocity $v$ as the relative velocity of the star-Earth system. This interpretation has been adopted by numerous textbook authors. As Moeller [13] states: ”This phenomenon, known as aberration, was observed by Bradley, who noticed that stars appear to undergo an annual collective motion in the sky. This apparent motion arises because the observed direction of a light ray from a star depends on the Earth’s velocity relative to the star.”
@@ 7.1 Inertial frame view of Observations of a Non-Inertial Observer
In this chapter, we introduce a novel approach to the theory of light aberration in non-inertial reference frames, offering a fresh perspective on this complex subject.
We begin by emphasizing a key principle: the laws of physics in an initial inertial frame must fully account for all physical phenomena, including those observed by non-inertial observers.
This principle was already applied in Chapter 5, where we analyzed phenomena in non-inertial (e.g., rotating) reference frames using the Langevin metric. This metric captures the electrodynamics of an accelerated light source from the viewpoint of measurements made by an accelerated observer—but interpreted from the perspective of an inertial observer.
Our findings there can be summarized as follows:
1. Inertial Frame Setup
We assume that the inertial frame has no history of acceleration and that spacetime, as perceived by an observer at rest, is described by the standard diagonal Minkowski metric: $ds^{2}=c^{2}dt^{2}-dx^{2}-dy^{2}-dz^{2}$. Applying a Galilean transformation $x\to x-vt$, $t\to t$ we obtain the metric of an accelerated emitter in the inertial frame, as presented in Eq. (11). In this coordinate system $(t,x,y,z)$, the devices of the inertial observer remain governed by the standard diagonal metric.
2. Non-Inertial Frame Transformation
To describe an accelerated observer with coordinates $(t_{n},x_{n},y_{n},z_{n})$, we apply the inverse Galilean transformation: $x\to x_{n}+vt$, with $t\to t_{n}$ unchanged. Substituting this into the Minkowski metric yields the Langevin metric, as shown in Eq. (15).
However, we argue that physical phenomena in accelerated frames can also be effectively analyzed within an inertial frame using the framework of standard Einsteinian special relativity and its well-established kinematic effects.
For instance, consider the Sagnac effect. From the perspective of an inertial laboratory frame, its interpretation becomes straightforward. The phase difference between counter-propagating waves arises from the relativistic velocity addition formula. That is, as viewed from the lab frame, the Sagnac effect emerges purely as a kinematic consequence of special relativity. This analysis assumes the speed of light in the lab frame is $c$, which—according to special relativity—is independent of the source’s motion when using Lorentz coordinates.
In contrast, analyzing the Sagnac effect within a rotating frame is more intricate. In such cases, authors typically employ a spacetime approach using a metric tensor—namely, the Langevin metric—within a four-dimensional Minkowski framework, to compute the propagation time differences between counter-propagating waves.
This leads us to a natural question: Can we analyze the aberration of light in the same way? Indeed, we propose applying the Langevin metric to describe light aberration in non-inertial frames as well.
That said, this approach requires a clock resynchronization procedure in the inertial frame. To maintain Lorentz coordinates for a moving light source, one must diagonalize the metric given in Eq. (11), as detailed in Section 3.5. The Lorentz coordinates can then be expressed in terms of the original coordinates $(t,x,y,z)$, using the transformation given in Eq. (13).
7.2 Relativity of Simultaneity
@@
It is particularly interesting to observe that geometric effects in our ordinary space are closely linked to the relativity of simultaneity. In Fig. 15, the transmitted light beam propagates at an angle of v/c-v/c, resulting in the phenomenon of light aberration. This naturally raises the question: relative to what does a light beam propagate within an accelerated frame when it undergoes an angular displacement of v/c-v/c? Consider an observer in the accelerated frame measuring the direction of the light beam. If the plane wavefront of the transmitted beam is focused by a lens, it will form a diffraction spot in the focal plane along the optical axis. Determining the direction of the optical axis with respect to the frame’s axes is equivalent to measuring the angular displacement. To detect the aberration of light within the accelerated frame, it is essential to establish a coordinate system with a well-defined reference direction. This necessitates a detailed examination of the method by which coordinates are assigned—a process that inherently involves a physical procedure. Ultimately, this method must provide a consistent framework for defining coordinates in both inertial and accelerated reference frames.
@@
Now, let us return to the observations of an accelerated observer as viewed from an inertial frame. In Chapter 3, we discussed how absolute time coordinatization can be transformed into Lorentz coordinatization. Within the inertial frame, this transformation can be understood as a change of the time variable according to $t\to t+xv/c^{2}$. The combined effect of the Galilean transformation and this variable change results in the Lorentz transformation when using absolute time coordinatization in the inertial frame. Consequently, in this framework, the speed of light remains independent of the motion of the source.
@@
From the preceding discussion, it follows that in an accelerated system, as viewed by an inertial observer, the electromagnetic axis assigned by the accelerated observer will not be parallel to the $z$-axis of the inertial frame (see Fig. 30). We conclude that, in the accelerated frame, the transmitted light beam propagates along the $z_{n}$-axis of the accelerated frame with an angular displacement of v/c-v/c.
@@
Now, we examine the angular displacement of the transmitted light relative to the reference beam within the accelerated frame. On one hand, the angular displacement of the transmitted light with respect to the reference beam from an emitter at rest in the accelerated frame is v/c-v/c. On the other hand, the angular displacement of the inertial reference light beam is 2v/c-2v/c (see Fig. 16). This implies that the accelerated observer can directly measure the angular displacement of the transmitted light with respect to the inertial frame’s reference light beam. Specifically, the relative angular displacement is $\theta_{a}(\mathrm{transmitted~beam})-\theta_{a}(\mathrm{inertial~reference~beam})=-v/c+2v/c=v/c$. This result is consistent with the measurements of the inertial observer, as expected  [^26] We have been discussing the electromagnetic axis. If we instead determine the reference directions in both frames using two relativistic electron sources, an important distinction arises. Unlike a reference light beam, a reference relativistic electron beam would not allow for an exact measurement of the transmitted light’s angular displacement within our experimental setup..
@@ 8 Kinematics of Wigner Rotation
The focus of this chapter is the Wigner rotation.  [^28] It is well known that the composition of non-collinear Lorentz boosts does not result in a simple boost but rather in a Lorentz transformation that includes both a boost and a spatial rotation. This rotation is known as the Wigner rotation [39, 40]. It is sometimes referred to as the Thomas rotation (see, e.g., [13, 41]). We previously introduced the Wigner rotation in the last chapter, where our discussion primarily centered on the aberration of light. However, the Wigner rotation is also closely related to the aberration of particles. As a relativistic kinematic effect, the Wigner rotation describes how the coordinate axes of a reference frame, moving along a curvilinear trajectory, rotate with respect to the axes of an inertial Lorentz frame. The primary objective of this chapter is to derive an expression for this effect. Throughout, we have aimed to minimize mathematical complexity while maintaining clarity.
@@ 8.1 Composition of Lorentz Boosts
Let us denote the three inertial frames by $K,K^{\prime},K$". The initial inertial frame is $K$, while $K^{\prime}$ is the rest frame of the particle, which has a velocity $\vec{v}=\vec{v}(\tau)$ relative to $K$, and KK" represents the particle’s rest frame at the next instant of proper time $\tau+d\tau$, moving with an infinitesimal velocity $d\vec{v^{\prime}}$ relative to $K^{\prime}$. All three inertial frames are assumed to be Lorentz frames. To clarify velocity measurements in different inertial frames: an observer in $K$ measures the velocity of frame $K^{\prime}$ as $\vec{v}$, while an observer in $K^{\prime}$ measures the velocity of frame KK" as $d\vec{v^{\prime}}$.
Now, let us examine the transformation of velocity in the theory of relativity. For a rectilinear motion along the $z$ axis, the velocity transformation follows the equation: $v_{z}(\tau+d\tau)=(dv^{\prime}_{z}+v_{z})/(1+v_{z}dv^{\prime}_{z}/c^{2})$. Just as with Galilean transformations, collinear Lorentz boosts commute—the order in which collinear boosts are applied does not affect the final result. This means that the composition of two Lorentz boosts in the same direction yields another Lorentz boost in that direction, with the total velocity given by the formula above.
However, when Lorentz boosts are applied in non-collinear directions, they no longer commute. While the composition of two Galilean boosts or two spatial rotations yields another transformation of the same kind (i.e., a boost or a rotation), the combination of two non-parallel Lorentz boosts results in a transformation that is not purely a boost—it includes a rotation known as the Wigner rotation.
To illustrate this, suppose frame $K^{\prime}$ moves with velocity $v_{z}$ along the $z$-axis relative to $K$ and in $K^{\prime}$, the particle experiences an infinitesimal acceleration in the $x^{\prime}$-direction, perpendicular to the trajectory of the initial inertial frame $K$. That is, frame KK" moves relative to $K^{\prime}$ with a small velocity $dv^{\prime}_{x}$ along the $x^{\prime}$-axis.
From the perspective of an inertial observer, time dilation occurs in the Lorentz frame moving with velocity $\vec{v}_{z}$ relative to the initial inertial frame, given by $dt/\gamma=d\tau$, where $\gamma=1/\sqrt{1-v_{z}^{2}/c^{2}}$. Due to this time dilation, the velocity increment in the initial inertial frame $dv_{x}$ corresponds to a velocity $dv^{\prime}_{x}=\gamma dv_{x}$ in $K^{\prime}$.
@@ 8.2 Wigner Rotation
Consider a succession of inertial frames $K\to K^{\prime}\to K$". From the perspective of the initial inertial frame, an observer in the frame KK" sees frame $K$ moving with velocity $v_{z}$ and velocity increment $\gamma d\vec{v}_{x}$. The corresponding rotation of the velocity direction in the KK" frame is given by $d\theta_{K"}=\gamma dv_{x}/v_{z}$ (Fig. 36). In the $K$ frame, the velocity rotation angle is $dv_{x}/v_{z}=d\theta$, with both rotations occurring in the same direction. If the proper and lab observers were to share a common three-dimensional space, this discrepancy would lead to a paradox — an apparent inconsistency in the theory. However, in Minkowski spacetime, no such paradox arises: the lab’s 3-space is rotated relative to the proper 3-space by an angle $d\Phi_{K"}=\gamma dv_{x}/v_{z}-dv_{x}/v_{z}=(\gamma-1)\d\theta$ in the same direction as the velocity vector’s rotation in the proper frame.
This angular difference $d\Phi_{K"}=(\gamma-1)\d\theta$ is the Wigner rotation — the rotation of the original frame’s axes as seen from the proper frame KK". In vector form, this can be expressed as $d\vec{\Phi}_{K"}=(\gamma-1)\vec{v}\times d\vec{v}/v^{2}$.
@@
Transforming this expression back to the lab frame $K$ yields $d\vec{\Phi}_{K}=(1-1/\gamma)\vec{v}\times d\vec{v}/v^{2}$ representing the Wigner rotation angle of the proper frame’s axes as observed in the lab frame. According to the principles of special relativity, this quantity must remain form-invariant under Lorentz transformations.  [^29] The correct expression for the Thomas (Wigner) rotation was first obtained by V. Ritus [42]. In deriving expressions for the Thomas (Wigner) rotation, the majority of authors (see e.g. [41]) was supposedly guided by the incorrect expression for Thomas (Wigner) rotation from Moeller’s monograph [13]. The expression obtained by Moeller is given by $\vec{\delta\Phi}=(1-\gamma)\vec{v}\times d\vec{v}/v^{2}=(1-\gamma)\vec{\delta\theta}$ (and subsequently expression for Thomas precession $\Omega_{\mathrm{T}}=(1-\gamma)\omega_{0}$). It should be noted that, in his monograph, Moeller stated several times that this expression is valid in the lab Lorentz frame. Clearly, this expression and correct result differ both in sign and magnitude. An analysis of why Moeller obtained an incorrect expression for the Thomas (Wigner) rotation in the lab frame is the focus of Ritus paper [43]. As analyzed in Ritus’s work [43], Moeller’s error is not computational but rather conceptual. The correct expression, derived decades earlier by several researchers, remained largely unnoticed amid the proliferation of inaccurate results, as discussed in the review [44]. We note that owing to the relativistic effect of time dilation in the reference frame that moves to the lab frame, the Wigner rotation angle in the proper frame is always $\gamma$ time higher than in the lab frame.  [^30] In 1986, M. Stranberg obtained an expression for the Thomas (Wigner) rotation correct both in the lab inertial frame and the comoving reference frame [45]. His paper is one of the few that explicitly states that the angle of Thomas (Wigner) rotation in the comoving frame is $\gamma$ times higher than in the lab frame [44].